### v1

@@ 1 Introduction
Existing RL methods for prescriptive maintenance can only optimize for a single objective, whereas most real-world maintenance optimization problems are multi-objective. Therefore, we introduce Multi-Objective Deep Centralized Multi-Agent Actor-Critic (MO-DCMAC), which is based on the DMCAC [4] and MOCAC [33] algorithms for maintenance planning and multi-objective reinforcement learning respectively. MO-DCMAC learns how to construct a maintenance plan in the same sequence as is currently done by asset managers, whereby the score is how well the maintenance plan performs according to the utility used.
If we consider infrastructural maintenance optimization as a multi-objective problem, we also need to consider what kind of scenario our method will be used. Hayes et al. [18] describe multiple scenarios for when a multi-objective approach is required. In most of these scenarios described by Hayes et al., the weighting of the objectives is not determined; however, asset managers who currently plan the maintenance know how to weigh these objectives with existing methodologies. One of these methodologies is the Failure Mode, Effects and Criticality Analysis (FMECA) methodology [7], where the utility (criticality score) of the different ways an asset can fail depends on both costs and risk (as well as other factors such as environment effects), which are combined in a non-linear manner. Therefore, according to Hayes et al. [18], we should consider the known utility function scenario because we know how to scalarize all the objectives into a singular value.
@@
Moreover, we formulate maintenance optimization for infrastructural assets as a multi-objective partially observable Markov decision process model (MOPOMDP) [39]. This MODPOMDP is formulated not only to handle uncertainty about the state of the components but also to model different reward signals that can be taken into account with a non-linear utility function. Additionally, we introduce a novel method for incorporating a probability value as a reward or objective. This reward value will ensure that episodic return is always a valid probability, which can be discounted. Lastly, we test MO-DCMAC on multiple environments where maintenance is planned for an infrastructural asset. For these environments, we use the historical quay walls of Amsterdam as a real-world use case (Sections 6.1.2 and 6.1.3).
@@ 4 Maintenance Optimization Problem
Before we explain our proposed method, MO-DCMAC, we first state the problem formulation of the maintenance optimization problem, for which we want to find an optimal policy and how this is formulated as Multi-Objective Partially Observable Markov decision process (MOPOMDP) [39]. In our explanation, we will also include how a MOPOMDP differs from a MOMDP and MDP.
@@ 4.1 Problem Formulation
An asset consists of $N$ components, whereby each component $\operatorname{c}_{i},i\in\{1,2,\dots,N\}$ can belong to one of $K\in\mathbb{N}$ component groups. The health state of each component $\operatorname{c}_{i}$ is represented by an integer variable $s_{t}^{i}\in\{1,2,\dots,|S|\}$, where $|S|$ is the number of possible health states of the component. Each component $\operatorname{c}_{i}$ has a current degradation rate $\operatorname{\tau}_{t}^{i}$ that can be increased until $\operatorname{\tau}_{\max}$. The degradation rate $\operatorname{\tau}_{t}^{i}$ dictates the probability that component $\operatorname{c}_{i}$ transitions from a healthy to a non-healthy state. Besides the degradation rate, to what component group $\operatorname{c}_{i}\in K$ belongs also influences the transition probability. This means some component groups might degrade faster than others, even if their degradation rate $\tau$ is higher.
@@
$\operatorname{\rho}_{t}=1-\prod_{k\in K}\prod^{\left|\operatorname{\mathcal{G}}^{k}\right|}_{i=1}\left(1-F{\left(\operatorname{\mathcal{G}}^{k}_{i}\right)}\right).$ (9)
@@ 4.2 Multi-Objective Partially Observable Markov Decision Process (MOPOMDP)
In our MOPOMDP formulation, we stated that the probability of collapse reward $\log{\left(1-\operatorname{\rho}_{t}\right)}$ as the inverse log probability of collapse occurring at timestep $t$. Using this inverse log probability is based on two key considerations. Firstly, using the log probability ensures that the episodic return of the probability of collapse $R_{\operatorname{\rho}}=\sum_{t=1}^{H}\log{\left(1-\operatorname{\rho}_{t}\right)}$ is a valid probability value. Secondly, $R_{\operatorname{\rho}}$ will be the log probability of no collapse occurring during the episode; therefore, we can use the base of the natural logarithm $e$ to obtain the probability of collapse over the whole episode easily:
@@ 5 Multi-Objective Deep Centralized Multi-Agent Actor-Critic (MO-DCMAC)
8:         Select actions $a_{t}=\left\{a^{\operatorname{c}_{1}}_{t},\dots,a^{\operatorname{c}_{N}}_{t},a^{\operatorname{\#}}_{t}\right\}\sim\pi_{\theta}(\cdot\mid\mathbf{b}_{t},\frac{t}{H},\vec{R}^{-}_{t})$.
@@
The first is the curse of dimensionality in the action space. The action space grows exponentially to the number of components and possible maintenance actions; for example, with 12 components and three possible actions, the action space would be $5314413^{12}=$ 531441. Andriotis and Papakonstantinou introduced Deep Centralized Multi-Agent Actor-Critic (DMCAC) [4], an actor-critic method, to alleviate this. DCMAC uses a factorized action space representation whereby the actor network has an output for each component that outputs an action for that component.
Secondly, our maintenance optimization problem is a multi-objective problem with a known non-linear utility function. Moreover, we want to learn a policy using the ESR criterion (Equation 7); however, this introduces significant complexity because if the utility function is non-linear, then the Bellman equation becomes invalidated (Equation 8) [18]. Reymond et al. introduced Multi-Objective Categorical Actor-Critic (MOCAC) [33] that utilizes a distributional critic, which learns a multi-variate distribution over the returns of the objective. This learned distribution is then used in combination with the accrued reward to directly optimize on the utility function.
Our proposed method, Multi-Objective Deep Centralized Multi-Agent Actor-Critic (MO-DCMAC), combines DMCAC [4] and MOCAC [33]. Combining these methods allows us to overcome the curse of dimensionality of action space by using the actor network of DCMAC and the ability to directly optimize over a non-linear utility function with the ESR criterion by using the critic network of MOCAC.
@@
The critic and most of the loss calculation in MO-DMCAC is based on MOCAC [33]. Reymond et al. introduce a distributional critic that learns the multi-variate distribution $Z$ of the return of the objectives. They based this on the work of Bellemare et al. [6], whereby the major difference is that Reymond et al. [33] learn a multi-variate distribution. The reason why the critic learns the distribution of the objectives, instead of the vectorial return, is that when the vectorial return is scalarized by a utility function $u$, it would learn under the SER criterion because the vectorial value-function learns the expected return (Equation 3). To ensure that MOCAC can learn an optimal policy under the ESR criterion, it requires the episodic return instead.
Therefore, the critic learns a distribution $Z$ through the following equation:
$V{\left(s_{t}\right)}=\sum_{i\in\left[\left\{1,\dots,\mathcal{N}\right\}\right]^{d}}\vec{z}_{i}Z_{\psi}{\left(\vec{z}_{i}\mid s_{t}\right)}.$ (12)
where $\vec{z}_{i}$ is the support atom for the vectorial return. Each atom $\vec{z}_{i}$ for $\forall i\in\left[\left\{1,\dots,\mathcal{N}\right\}\right]^{d}$ is calculated by,
@@
Due to Equations 12 and 13, the critic learns the discretized distribution of the objectives instead of the expected return. Because $\vec{z}$ does not represent the expected return but is defined as a return, MOCAC [33] can learn under the ESR criterion. Therefore, the critic can be used, in combination with the accrued reward $\vec{R}^{-}_{t}=\sum_{k=0}^{t-1}\gamma^{k}\vec{r}_{k}$ to calculate a preference score $u_{t}$ under ESR criterion by:
$u_{t}=\sum_{i\in\left[\left\{1,\dots,\mathcal{N}\right\}\right]^{d}}u{\left(\vec{R}^{-}_{t}+\vec{z}_{i}\right)}Z_{\psi}{\left(\vec{z}_{i}\mid s_{t}\right)}.$ (14)
With the utility score being calculated through Equation 14, the advantage $A(s_{t},a_{t})$ of MOCAC [33] and MO-DCMAC is calculated through:
$A{\left(s_{t},a_{t}\right)}=$ $\Bigg{(}\sum_{i\in\left[\left\{1,\dots,\mathcal{N}\right\}\right]^{d}}u\left(\vec{R}^{-}_{t}+\gamma^{t}\left(\vec{r}_{t}+\gamma\vec{z}_{i}\right)\right)$
@@
$-\Bigg{(}\sum_{i\in\left[\left\{1,\dots,\mathcal{N}\right\}\right]^{d}}u\left(\vec{R}^{-}_{t}+\gamma^{t}\vec{z}_{i}\right)$
@@
Reymond et al. did not normalize the advantage in MOCAC. However, we found that normalizing significantly increased training stability; therefore, the final advantage used is:
$\hat{A}{(a_{t},s_{t})}=\frac{A{(a_{t},s_{t})}-\mu_{A}}{\sigma_{A}}$ (16)
@@
$\cdot Z_{\psi}{\left(\vec{z}_{i}\mid s_{t+1}\right)}\Bigg{)},\forall i\in\left[\left\{1,\dots,\mathcal{N}\right\}\right]^{d},$ (17)
@@
$\mathcal{L}^{\pi}{(\pi)}=-w_{t}\left(\sum_{j=1}^{N}\log{\left(\pi^{\operatorname{c}_{j}}_{\theta}{\left(a^{\operatorname{c}_{j}}_{t}\mid\mathbf{b}_{t}\right)}\right)}\right)A{\left(\mathbf{b}_{t},a_{t}\right)},$ (19)
where $w_{t}=\min\left(q,\frac{\pi\left(a_{t}\mid s_{t}\right)}{\mu\left(a_{t}\mid s_{t}\right)}\right)$ is a truncated importance sampling [42] to address the high variance that log summation may introduce and $\mathbf{b}_{t}$ is the belief at timestep $t$ because DCMAC is used to find an optimal policy in a POMDP. The reason a truncated importance sampling is needed is that DCMAC is an off-policy method because it collects training examples in a replay buffer and samples during training from it. In contrast, MO-DCMAC is an on-policy method, meaning we do not reuse training examples similar to MOCAC [33]. Therefore, the importance sampling is not necessary anymore due to $\frac{\pi\left(a_{t}\mid s_{t}\right)}{\pi\left(a_{t}\mid s_{t}\right)}=1$. Additionally, we introduce another separate head that selects a global action head from $A^{\operatorname{\#}}$. Therefore, the modified policy loss calculation will be:
$\mathcal{L}^{\pi}{(\pi)}=-\bigg{(}$ $\sum_{j=1}^{N}\log{\left(\pi^{\operatorname{c}_{j}}_{\theta}{\left(a^{\operatorname{c}_{j}}_{t}\mid\mathbf{b}_{t}\right)}\right)}$
@@
where $\mathcal{H}(\pi)$ is the entropy bonus from $\pi$.
@@ 6.2 Utility Functions
As previously discussed, the input for the utility functions is the episodic return of cost $R_{\operatorname{\mathcal{C}}}=\sum_{t=0}^{H}\gamma^{t}-\operatorname{\mathcal{C}}_{t}$ and the probability of collapse $R_{\operatorname{\rho}}=\sum_{t=0}^{H}\gamma^{t}\log(1-\operatorname{\rho}_{t})$, where the probability of collapse is calculated to $P_{\operatorname{\rho}}=1-e^{R_{\operatorname{\rho}}}$ (Equation 11). This probability value $P_{\operatorname{\rho}}$ is the probability that during the planned period, a collapse can occur.
@@ 6.3 Baseline Policies
We did consider other methods; however, they are highly unlikely to result in a more optimal maintenance policy. Other reinforcement learning methods, such as DCMAC [4], cannot learn an optimal policy, even with adjustments for learning a multi-objective optimization. Reymond et al. [33] showed in their comparison of MOCAC to other baseline methods that MOCAC consistently learns a better policy. Moreover, other methods for utility-based maintenance optimization, such as genetic algorithms [8, 11, 15], cannot scale to problem instances of this size.

### v2

@@ 1 Introduction
Existing RL methods for prescriptive maintenance can only optimize for a single objective, whereas most real-world maintenance optimization problems are multi-objective. Therefore, we introduce Multi-Objective Deep Centralized Multi-Agent Actor-Critic (MO-DCMAC) for multi-objective prescriptive maintenance planning. MO-DCMAC is based on the DCMAC [4] and MOCAC [33] algorithms for maintenance planning and multi-objective reinforcement learning, respectively. MO-DCMAC learns how to construct a maintenance plan in the same sequence as is currently done by asset managers, whereby the score is how well the maintenance plan performs according to the utility used.
If we consider infrastructural maintenance optimization as a multi-objective problem, we also need to consider what kind of scenario our method will be used. Hayes et al. [18] describe multiple scenarios for when a multi-objective approach is required. In most of these scenarios described by Hayes et al., the weighting of the objectives is not determined; however, asset managers who currently plan the maintenance know how to weigh these objectives with existing methodologies. One of these methodologies is the Failure Mode, Effects, and Criticality Analysis (FMECA) [7], which evaluates the utility (criticality score) of various asset failure modes based on a combination of costs, risks, and other factors such as environmental effects, all integrated in a non-linear manner. Therefore, according to Hayes et al. [18], we should consider the known utility function scenario because we know how to scalarize all the objectives into a singular value.
@@
Moreover, we formulate maintenance optimization for infrastructural assets as a multi-objective partially observable Markov decision process model (MOPOMDP) [39]. This MOPOMDP is formulated not only to handle uncertainty about the state of the components but also to model different reward signals that can be taken into account with a non-linear utility function. Additionally, we introduce a novel method for incorporating a probability value as a reward or objective. This reward value will ensure that episodic return is always a valid probability, which can be discounted. Lastly, we test MO-DCMAC on multiple environments where maintenance is planned for an infrastructural asset. For these environments, we use the historical quay walls of Amsterdam as a real-world use case (Sections 6.1.2 and 6.1.3).
@@ 4 Maintenance Optimization Problem
Before we explain our proposed method, MO-DCMAC, we first state the problem formulation of the maintenance optimization problem, for which we want to find an optimal policy and how this is formulated as a Multi-Objective Partially Observable Markov decision process (MOPOMDP) [39]. In our explanation, we will also include how a MOPOMDP differs from a MOMDP and MDP.
@@ 4.1 Problem Formulation
An asset consists of $N$ components, whereby each component $\operatorname{c}_{i},i\in\{0,1,\dots,N-1\}$ can belong to one of $K\in\mathbb{N}$ component groups. The health state of each component $\operatorname{c}_{i}$ is represented by an integer variable $s_{t}^{i}\in\{0,1,\dots,|S|-1\}$, where $|S|$ is the number of possible health states of the component. Each component $\operatorname{c}_{i}$ has a current degradation rate $\operatorname{\tau}_{t}^{i}$ that can be increased until $\operatorname{\tau}_{\max}$. The degradation rate $\operatorname{\tau}_{t}^{i}$ dictates the probability that component $\operatorname{c}_{i}$ transitions from a healthy to a non-healthy state. Besides the degradation rate, to what component group $\operatorname{c}_{i}\in K$ belongs also influences the transition probability. This means some component groups might degrade faster than others, even if their degradation rate $\tau$ is higher.
@@
$\operatorname{\rho}_{t}=1-\prod_{k\in K}\prod^{\left|\operatorname{\mathcal{G}}^{k}\right|-1}_{i=0}\left(1-F{\left(\operatorname{\mathcal{G}}^{k}_{i}\right)}\right).$ (9)
@@ 4.2 Multi-Objective Partially Observable Markov Decision Process (MOPOMDP)
In our MOPOMDP formulation, we stated that the probability of collapse reward $\log{\left(1-\operatorname{\rho}_{t}\right)}$ as the inverse log probability of collapse occurring at timestep $t$. Using this inverse log probability is based on two key considerations. Firstly, using the log probability ensures that the episodic return of the probability of collapse $R_{\operatorname{\rho}}=\sum_{t=0}^{H}\log{\left(1-\operatorname{\rho}_{t}\right)}$ is a valid probability value. Secondly, $R_{\operatorname{\rho}}$ will be the log probability of no collapse occurring during the episode; therefore, we can use the base of the natural logarithm $e$ to obtain the probability of collapse over the whole episode easily:
@@ 5 Multi-Objective Deep Centralized Multi-Agent Actor-Critic (MO-DCMAC)
8:         Select actions $a_{t}=\left\{a^{\operatorname{c}_{0}}_{t},\dots,a^{\operatorname{c}_{N-1}}_{t},a^{\operatorname{\#}}_{t}\right\}\sim\pi_{\theta}(\cdot\mid\mathbf{b}_{t},\frac{t}{H},\vec{R}^{-}_{t})$.
@@
The first is the curse of dimensionality in the action space. The action space grows exponentially to the number of components and possible maintenance actions; for example, with 12 components and three possible actions, the action space would be $5314413^{12}=$ 531441. Andriotis and Papakonstantinou introduced Deep Centralized Multi-Agent Actor-Critic (DCMAC) [4], an actor-critic method, to alleviate this. DCMAC uses a factorized action space representation whereby the actor network has an output for each component that outputs an action for that component.
Secondly, our maintenance optimization problem is a multi-objective problem with a known non-linear utility function. Moreover, we want to learn a policy using the ESR criterion (Equation 7); however, this introduces significant complexity because if the utility function is non-linear, then the Bellman equation becomes invalidated (Equation 8) [18]. Reymond et al. introduced Multi-Objective Categorical Actor-Critic (MOCAC) [33] that utilizes a distributional critic, which learns a multi-variate distribution over the returns of the objective. MO-DCMAC requires this multi-variate distribution over the return, introduced by Reymond et al. [33], to be able to learn an optimal policy with the ESR criterion in a multi-objective problem setting. We explain this further in this section.
Our proposed method, Multi-Objective Deep Centralized Multi-Agent Actor-Critic (MO-DCMAC), combines DCMAC [4] and MOCAC [33]. Combining these methods allows us to overcome the curse of dimensionality of action space by using the actor network of DCMAC and the ability to directly optimize over a non-linear utility function with the ESR criterion by using the critic network of MOCAC.
@@
The critic and most of the loss calculations in MO-DCMAC are based on previous work introduced by Reymond et al. [33]. Reymond et al. introduce a distributional critic that learns the multi-variate distribution $Z$ of the return of the objectives. They based this on the work of Bellemare et al. [6], whereby the major difference is that Reymond et al. [33] learn a multi-variate distribution. The reason why the critic learns the distribution of the objectives instead of the expected vectorial return is that when the expected vectorial return is scalarized by a utility function $u$, it would learn under the SER criterion because the vectorial value-function learns the expected return (Equation 3). To ensure that MO-DCMAC can learn an optimal policy under the ESR criterion, episodic return is required instead [33].
Therefore, we utilize the critic introduced by Reymond et al. [33] to learn a distribution $Z$ through the following equation:
$V{\left(s_{t}\right)}=\sum_{i\in\left[\left\{0,\dots,\mathcal{N}-1\right\}\right]^{d}}\vec{z}_{i}Z_{\psi}{\left(\vec{z}_{i}\mid s_{t}\right)}.$ (12)
where $\vec{z}_{i}$ is the support atom for the vectorial return. Each atom $\vec{z}_{i}$ for $\forall i\in\left[\left\{0,\dots,\mathcal{N}-1\right\}\right]^{d}$ is calculated by,
@@
Due to Equations 12 and 13 introduced by Reymond et al. [33], the critic learns the discretized distribution of the objectives instead of the expected return. Because $\vec{z}$ does not represent the expected return but is defined as a return, MO-DCMAC can learn under the ESR criterion. Therefore, in combination with the accrued reward $\vec{R}^{-}_{t}=\sum_{k=0}^{t-1}\gamma^{k}\vec{r}_{k}$, we can calculate a preference score $u_{t}$ [33] under the ESR criterion by:
$u_{t}=\sum_{i\in\left[\left\{0,\dots,\mathcal{N}-1\right\}\right]^{d}}u{\left(\vec{R}^{-}_{t}+\vec{z}_{i}\right)}Z_{\psi}{\left(\vec{z}_{i}\mid s_{t}\right)}.$ (14)
With the preference score $u_{t}$ being calculated through Equation 14, the advantage $A(s_{t},a_{t})$ of MO-DCMAC is calculated through Equation 15 [33]
$A{\left(s_{t},a_{t}\right)}=$ $\Bigg{(}\sum_{i\in\left[\left\{0,\dots,\mathcal{N}-1\right\}\right]^{d}}u\left(\vec{R}^{-}_{t}+\gamma^{t}\left(\vec{r}_{t}+\gamma\vec{z}_{i}\right)\right)$
@@
$-\Bigg{(}\sum_{i\in\left[\left\{0,\dots,\mathcal{N}-1\right\}\right]^{d}}u\left(\vec{R}^{-}_{t}+\gamma^{t}\vec{z}_{i}\right)$
@@
Reymond et al. [33] did not originally normalize the advantage. However, we found that normalizing significantly increased training stability; therefore, the final advantage used is:
$\hat{A}{(a_{t},s_{t})}=\frac{A{(a_{t},s_{t})}-\mu_{A}}{\sigma_{A}},$ (16)
where $\mu_{A}$ is the mean advantage of the current batch and $\sigma_{A}$ is the standard deviation of the advantage of the current batch.
@@
$\cdot Z_{\psi}{\left(\vec{z}_{i}\mid s_{t+1}\right)}\Bigg{)},$
$\forall i\in\left[\left\{0,\dots,\mathcal{N}-1\right\}\right]^{d},$ (17)
@@
$\mathcal{L}^{\pi}{(\pi)}=-w_{t}\left(\sum_{j=0}^{N-1}\log{\left(\pi^{\operatorname{c}_{j}}_{\theta}{\left(a^{\operatorname{c}_{j}}_{t}\mid\mathbf{b}_{t}\right)}\right)}\right)A{\left(\mathbf{b}_{t},a_{t}\right)},$ (19)
where $w_{t}=\min\left(q,\frac{\pi\left(a_{t}\mid s_{t}\right)}{\mu\left(a_{t}\mid s_{t}\right)}\right)$ is a truncated importance sampling [42] to address the high variance that log summation may introduce, $c_{j}$ the component for which an action is selected, and $\mathbf{b}_{t}$ is the belief at timestep $t$ because DCMAC is used to find an optimal policy in a POMDP. The reason a truncated importance sampling is needed is that DCMAC is an off-policy method because it collects training examples in a replay buffer and samples during training from it. In contrast, MO-DCMAC is an on-policy method, meaning we do not reuse training examples similar to MOCAC [33]. Therefore, the importance sampling is not necessary anymore due to $\frac{\pi\left(a_{t}\mid s_{t}\right)}{\pi\left(a_{t}\mid s_{t}\right)}=1$. Additionally, we introduce another separate head that selects a global action head from $A^{\operatorname{\#}}$. Therefore, the modified policy loss calculation will be:
$\mathcal{L}^{\pi}{(\pi)}=-\bigg{(}$ $\sum_{j=0}^{N-1}\log{\left(\pi^{\operatorname{c}_{j}}_{\theta}{\left(a^{\operatorname{c}_{j}}_{t}\mid\mathbf{b}_{t}\right)}\right)}$
@@
where $\mathcal{H}(\pi)$ is the entropy bonus from $\pi$, $\lambda_{\text{val}}$ the coefficient for the value loss and $\lambda_{\text{ent}}$ the coefficient for the entropy bonus. Both $\lambda_{\text{value}}$ and $\lambda_{\text{ent}}$ are hyperparameters and control the contribution of the value and entropy loss, respectively.
@@ 6.2 Utility Functions
As previously discussed, the input for the utility functions is the episodic return of cost $R_{\operatorname{\mathcal{C}}}=\sum_{t=0}^{H}\gamma^{t}(-\operatorname{\mathcal{C}}_{t})$ and the log probability of collapse $R_{\operatorname{\rho}}=\sum_{t=0}^{H}\gamma^{t}\log(1-\operatorname{\rho}_{t})$, where the log probability of collapse $R_{\operatorname{\rho}}$ is calculated to $P_{\operatorname{\rho}}$ (Equation 11). This probability value $P_{\operatorname{\rho}}$ is the probability that during the planned period, a collapse can occur.
@@ 6.3 Baseline Policies
We did not include other single-objective RL methods or multi-objective RL methods since they cannot learn optimal policies for the given problem setting. Reymond et al. [33] showed that single-objective RL methods trained on only the utility score, given at the last state, result in highly unstable policies that always underperform. Other multi-objective RL methods cannot learn with the ESR criterion [18], and therefore, we cannot compare the results of MO-DCMAC with the results of these methods. We did not include MOCAC [33] since it would require substantial adjustments for it to be able to function in our problem setting. Lastly, other methods for utility-based maintenance optimization, such as genetic algorithms [8, 11, 15], cannot scale to problem instances of this size.